An ideal step-down transformer has a primary coil of 350 turns and a secondary coil of 10 turns. Its primary coil is plugged into an outlet with 220 V (AC). If the current in the secondary coil is 14.35 A, what current does the primary coil draw?
Answer #1

Relation between current and turns of the primary and secondary in the transformer is

I1/I2=N2/N1

I1=(N2/N1)*I2

=(10/350)*14.35 A

=0.41 A